Determine the dimension and hence the basis for the vector space spanned by the vectors (-1,1,3),(2,3,4),(3,0,-5) and (-2,1,0)


1
Expert's answer
2021-08-16T15:25:17-0400
"\begin{pmatrix}
 -1 & 2 & 3 & -2 \\
 1 & 3 & 0 & 1 \\
 3 & 4 & -5 & 0 \\
\end{pmatrix}"

"R_1=-R_1"


"\begin{pmatrix}
 1 & -2 & -3 & 2 \\
 1 & 3 & 0 & 1 \\
 3 & 4 & -5 & 0 \\
\end{pmatrix}"

"R_2=R_2-R_1"


"\begin{pmatrix}
 1 & -2 & -3 & 2 \\
 0 & 5 & 3 & -1 \\
 3 & 4 & -5 & 0 \\
\end{pmatrix}"

"R_3=R_3-3R_1"


"\begin{pmatrix}
 1 & -2 & -3 & 2 \\
 0 & 5 & 3 & -1 \\
 0 & 10 & 4 & -6 \\
\end{pmatrix}"

"R_2=R_2\/5"


"\begin{pmatrix}
 1 & -2 & -3 & 2 \\
 0 & 1 & 3/5 & -1/5 \\
 0 & 10 & 4 & -6 \\
\end{pmatrix}"

"R_1=R_1+2R_2"


"\begin{pmatrix}
 1 & 0 & -9/5 & 8/5 \\
 0 & 1 & 3/5 & -1/5 \\
 0 & 10 & 4 & -6 \\
\end{pmatrix}"

"R_3=R_3-10R_2"


"\begin{pmatrix}
 1 & 0 & -9/5 & 8/5 \\
 0 & 1 & 3/5 & -1/5 \\
 0 & 0 & -2 & -4 \\
\end{pmatrix}"

"R_3=R_3\/(-2)"


"\begin{pmatrix}
 1 & 0 & -9/5 & 8/5 \\
 0 & 1 & 3/5 & -1/5 \\
 0 & 0 & 1 & 2 \\
\end{pmatrix}"

"R_1=R_1+9R_3\/5"


"\begin{pmatrix}
 1 & 0 & 0 & 26/5 \\
 0 & 1 & 3/5 & -1/5 \\
 0 & 0 & 1 & 2 \\
\end{pmatrix}"

"R_2=R_2-3R_3\/5"

"\begin{pmatrix}
 1 & 0 & 0 & 26/5 \\
 0 & 1 & 0 & -7/5 \\
 0 & 0 & 1 & 2 \\
\end{pmatrix}"

"rank A=3=>" the dimension is 3.


the basis for the vector space is "\begin{bmatrix}
 -1 \\
 1 \\
3
\end{bmatrix}, \begin{bmatrix}
 2 \\
 3 \\
4
\end{bmatrix},\begin{bmatrix}
 3 \\
 0 \\
-5
\end{bmatrix}"